-8.314 T. In XE = 8.4 (1500-T) -8.314T, In(1-X)= 8.4 (1300-TE)

Find Xe and Te
Answer #1

(11 – 8-314 Te In Xh = 8:4 (1500-TE) -8.314 Te In (1-x) = 8.4 (1300-TE). we know, In (2) = )= x - 2 + 1 - 27 In (1-2)=-[n- 2E 7 in In x A 840 11639.6 = -0.072167 E of 0.930375 so, I TE z 1400 Xe > 0-230315